OpenStudy (anonymous):

a coin is biased so that the probability of obtaining a head is 3/7. if the coin is tossed 3 times the probability of obtaining exactly 2 heads is ????

OpenStudy (anonymous):

binomial distribution for this one 2 heads, one tails probability of head is \(\frac{3}{7}\) so probability of tails is what?

OpenStudy (anonymous):

4/7

OpenStudy (anonymous):

ok good so heads twice \(\frac{3}{7}\times \frac{3}{7}\)

OpenStudy (anonymous):

then tails once \(\frac{3}{7}\times \frac{3}{7}\times \frac{4}{7}\)

OpenStudy (anonymous):

and there are three different ways to get this h, h , t h, t, h t, h, h so multiply the above number by 3

OpenStudy (anonymous):

i.e. final answer is \[P(x=2)=3\times \left(\frac{3}{7}\right)^2\times \frac{4}{7}\]

OpenStudy (anonymous):

general formula is \[P(x=k)=\binom{n}{k}p^k(1-p)^{n-k}\]

OpenStudy (anonymous):

in your example \(n=3,k=2, \binom{3}{2}=3, p=\frac{3}{7}, 1-p=\frac{4}{7}\)

OpenStudy (anonymous):

I'm writing these down. They are very helpful

OpenStudy (anonymous):

ok let me know if you have any questions about the numbers or the notation

OpenStudy (anonymous):

If the question asked 4 tosses then that will mean two heads and two tails

OpenStudy (anonymous):

yes so it would be \(P(x=2)=\binom{4}{2}\left(\frac{3}{7}\right)^2\left(\frac{4}{7}\right)^2\)

OpenStudy (anonymous):

do you know the notation \(\binom{4}{2}\) ? you might have seen it written as \(_4C_2\)

OpenStudy (anonymous):

\(\binom{n}{k}\) reads "n choose k"

OpenStudy (anonymous):

4 is the number of tosses

OpenStudy (anonymous):

The 2 is for the number of heads

OpenStudy (anonymous):

Am I right?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

suppose instead you tossed this unfair coin 10 times and you wanted the probability you got exactly 3 heads could you find it?

OpenStudy (anonymous):

Ill try give me some time and ill be right back with you :)

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

can we please go through it together?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

i guess n=10 and k=3

OpenStudy (anonymous):

?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

n= number of times the coin is tossed

OpenStudy (anonymous):

right

OpenStudy (anonymous):

k= number of heads which is 3

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but you don't have to think of it as a "coin" it can be any experiment where there are only two possible outcomes a coin is just a good model, because it can only be "heads" or "tails" but it could be "rain" or "shine" or " yes" or "no" usually it is denoted by "success" or "failure"

OpenStudy (anonymous):

but i interrupt you are doing fine, keep going

OpenStudy (anonymous):

P(x=3) = (10 3) (3/7)^3 (1-3/7)^10-3

OpenStudy (anonymous):

yes, exactly

OpenStudy (anonymous):

however... do not get married to the formula you don't want to be looking at \(1-\frac{3}{7}\) but rather \(\frac{4}{7}\)

OpenStudy (anonymous):

also of course \(10-3=7\)

OpenStudy (anonymous):

what do u mean?

OpenStudy (anonymous):

i mean you want to know \(1-p\) and \(n-k\) before you put it in to the formula

OpenStudy (anonymous):

okay now i get it

OpenStudy (anonymous):

so while you are completely correct, you really want to write \[P(x=3)=\binom{10}{3}\left(\frac{3}{7}\right)^3\ \left(\frac{4}{7}\right)^7\]

OpenStudy (anonymous):

now it makes complete sense 3 head 7 tails \(\binom{10}{3}\) ways to get it

OpenStudy (anonymous):

question is, do you know how to compute \(\binom{10}{3}\) ?

OpenStudy (anonymous):

mmmm im not sure

OpenStudy (anonymous):

for large numbers you would use a calculator with that feature it means the number of ways to choose 3 from a set of 10 here \[\binom{10}{3}=\frac{10\times 9\times 8}{3\times 2}=10\times 3\times 4=120\]

OpenStudy (anonymous):

i didnt get it

OpenStudy (anonymous):

you mean you don't understand how i got the answer right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok lets do a smaller example lets find \(\binom{6}{2}\)

OpenStudy (anonymous):

sometimes written as \(_6C_2\)

OpenStudy (anonymous):

the number of ways to choose 2 items from a set of 6 you have 6 choices for the first item then 5 for the second by the counting principle the number of ways to do this together is \(6\times 5\) but order does not matter in choosing, so choosing \(a\) and then \(b\) is no different from choosing \(b\) and then \(a\) so \(6\times 5\) overcounts by a factor of \(2\) therefore \[\binom{6}{2}=\frac{6\times 5}{2}=3\times 5=15\]

OpenStudy (anonymous):

you can find these numbers also in pascal's triangle

OpenStudy (anonymous):

so (3 2) = 3x 2/ 2

OpenStudy (anonymous):

now lets try \[\binom{10}{3}\] 10 choices for the first item 9 for the second 8 for the third so by counting principle \(10\times 9\times 8\) but this over counts by the number of permutations of 3 items, which is \(3!=3\times 2\) so \[\binom{10}{3}=\frac{10\times 9\times 8}{3\times 2}\]

OpenStudy (anonymous):

yes you are right

OpenStudy (anonymous):

\(\binom{3}{2}=\frac{3\times 2}{2}=3\)

OpenStudy (anonymous):

what is this rule called?

OpenStudy (anonymous):

wouldn't really use the formula in that case because it is more or less clear that the number of ways to choose 2 from a set of 3 is the same as the number of ways to choose one from a set of 3 which is 3

OpenStudy (anonymous):

\[\binom{n}{k}\] is read "n choose k"

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

there is a formula for it as well

OpenStudy (anonymous):

so how ccan we find the final answer?

OpenStudy (anonymous):

\[\binom{n}{k}=\frac{n!}{k!(n-k)!}\] but just like the formula we had above, you should not get married to it the denominator tells you what to cancel

OpenStudy (anonymous):

the answer for the original question, or the answer for \(\binom{10}{3}\) ?

OpenStudy (anonymous):

oh now i see that is the binomial coefficent

OpenStudy (anonymous):

for (10 3)

OpenStudy (anonymous):

yes, it is the binomial coefficient

OpenStudy (anonymous):

\[\binom{10}{3}=\frac{10\times 9\times 8}{3\times 2}=10\times 3\times 4=120\]

OpenStudy (anonymous):

is that the final answer or do we have to work out (3/7)^3 (4/7)^7

OpenStudy (anonymous):

yes but with a calculator i would not try this by hand

OpenStudy (anonymous):

i got 0.00156...

OpenStudy (anonymous):

your answer times 120

OpenStudy (anonymous):

man you rock. thanks for the help. i really appreciate it man seriously

OpenStudy (anonymous):

you are quite welcome hope it makes some sense, and hope you ace your probability class

OpenStudy (anonymous):

i hope so. yes i wouldnt have understood it if it wasnt for you :)